# Feature-aware Hypergraph Generation via Next-Scale Prediction

### v1

@@ 3.3 Budgeted Expansion and Refinement
During expansion, since our framework only conditions on the total number of nodes in the final hypergraph, we maintain a single node budget vector $\mathbf{b}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ for the left-side (node) partition and discard the right-side (hyperedge) budget. Each cluster is split into multiple child nodes, which initially inherit both the parent’s budget and feature vector. A subsequent refinement step then determines: i) how the parent’s budget should be divided among its children, and ii) how to update the children’s feature vectors.
More specifically, the expansion step duplicates vertices and hyperedges, with child nodes inheriting all of their parent’s connections. The refinement step then i) removes edges that should not persist at the finer resolution, ii) redistributes the integer budget among the child nodes according to a probability-like split vector, and iii) replaces the coarse features with new refined predictions.
@@
Given a bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$ and two cluster size vectors $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$, $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$, denoting the expansion size of each node, let $\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b}^{\text{expanded}},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$ denote the expansion of $B$, whose node sets, budgets, and features are given by:
$\begin{split}&\tilde{\mathcal{V}}_{L}=\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(|\mathcal{V}_{L}|)},\text{~{}where~{}~{}}\mathcal{V}_{L}^{(p)}=\{v_{L}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{L}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{L}|,\\ &\tilde{\mathcal{V}}_{R}=\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(|\mathcal{V}_{R}|)},\text{~{}where~{}~{}}\mathcal{V}_{R}^{(p)}=\{v_{R}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{R}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{R}|,\\ &\left(\mathbf{b}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\mathbf{b}_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{L}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\left(\mathbf{F}_{L}\right)_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{R}^{\text{expanded}}\right)_{v_{R}^{(p,i)}}=\left(\mathbf{F}_{R}\right)_{\mathcal{V}_{R}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{R}[p],1\leq p\leq|\mathcal{V}_{R}|.\end{split}$ (3)
@@
Expansion thus acts as a clone-and-rewire operation: vertices and hyperedges are duplicated, and each child inherits every incident edge of its parent.
@@
Given a bipartite graph $\tilde{B}=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$, an edge selection vector $\mathbf{e}\in\{0,1\}^{|\mathcal{E}|}$, a budget split vector $\mathbf{f}\in[0,1]^{|\tilde{\mathcal{V}}_{L}|}$, satisfying $\sum_{v\in\mathcal{V}_{L}^{(p)}}\mathbf{f}_{v}=1$ for all cluster $\mathcal{V}_{L}^{(p)}$ in $\mathcal{V}_{L}$, and two feature refinement vectors $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ with the same dimensions as $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$, let $B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\mathcal{E},\lceil\mathbf{b}\odot\mathbf{f}\rfloor,\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ denote the refinement of $\tilde{B}$, where $\mathcal{E}\subseteq\tilde{\mathcal{E}}$ such that the $i$-th edge $e_{(i)}\in\mathcal{E}$ if and only if $\mathbf{e}[i]=1$.
@@
The generation of a hypergraph with $N$ nodes proceeds in three main stages:
@@
    Initialization. Start from a minimal bipartite graph: $B^{(L)}=(\{1\},\{2\},\{(1,2)\})$, consisting of a single node on each side connected by one edge. The left-side node is assigned the full node budget, i.e., $\mathbf{b}=(N)$. If node and hyperedge features need to be generated, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$ are initialized as zero matrices.
@@
[…]
@@
3.4 Probabilistic Modeling
We now present a formalization of our learning framework, generalizing [11, 6]. Let $\{H^{(1)},\ldots,H^{(N)}\}$ denote a set of i.i.d. hypergraph instances. Our aim is to approximate the unknown generative process by learning a distribution $p(H)$. We model the marginal likelihood of each hypergraph $H$ as a sum over the likelihoods of its bipartite representation’s expansion sequences:
$p(H)=p(B)=\sum_{\varpi\in\Pi(B)}p(\varpi),$ (4)
where $\Pi(B)$ denotes the set of valid expansion sequences from a minimal bipartite graph to the full bipartite representation $B$ corresponding to $H$. Each intermediate $B^{(l-1)}$ is generated by refining its predecessor, in accordance with Definitions 4 and 6.
Assuming a Markovian generative structure, the likelihood of a specific expansion sequence $\varpi$ is factorized as:
$p(\varpi)=\underbrace{p(B^{(L)})}_{1}\prod_{l=L}^{1}p(B^{(l-1)}|B^{(l)})=\prod_{l=L}^{1}p(\mathbf{e}^{(l-1)}|\tilde{B}^{(l-1)})p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (5)
To simplify the modeling process and avoid learning two separate distributions $p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})$ and $p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})$, we rearrange terms as follows:
$p(\varpi)=p(\mathbf{e}^{0}|\tilde{B}^{0})p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})\left[\prod_{l=L-1}^{1}p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})\right],$ (6)
where $p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)})=p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)}|B^{(L)})$.
We assume that the variables $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{b}^{(l)}$, $\mathbf{F}_{L}^{(l)}$, and $\mathbf{F}_{R}^{(l)}$ are conditionally independent of $\tilde{B}^{(l)}$ when conditioned on $B^{(l)}$:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)},\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (7)
This allows us to write the combined likelihood as:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{e}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (8)
3.5 Flow-matching
We use a flow-matching framework [21] with an endpoint parameterization similar to [22], or alternatively a denoising diffusion model [23] with linear interpolation between the prior and target distribution. The objective is to align two probability distributions $p_{0}$ and $p_{1}$ by evolving the points from the initial distribution $p_{0}$ to the target distribution $p_{1}$. This is achieved by defining a flow through a time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the following ordinary differential equation (ODE) $\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\mathbf{x}_{0}\sim p_{0}$, where $\mathbf{f}(\mathbf{x},t)$ is learned such that, after integrating the ODE, samples from $p_{1}$ can be generated.
With endpoint parametrization, the model learns to predict the endpoint $\mathbf{x}_{1}$ of the trajectory. The flow $\mathbf{f}(\mathbf{x},t)$ can then be recovered from the predicted endpoint $\hat{\mathbf{x}}_{1}(\mathbf{x},1)$ using the relation:
$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x},1)-\mathbf{x}_{t}}{1-t}.$ (9)
To train the model, the loss function minimizes the squared difference between the predicted endpoint $\hat{\mathbf{x}}_{1}(x_{t},t)$ and the actual endpoint $\mathbf{x}_{1}$:
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left(\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right)^{2}\right],\text{~{}~{}where~{}~{}}\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}.$ (10)
For the expansion process and edge selection, we set $p_{0}$ as the Gaussian distribution. For budget splits, as their support is constrained to the simplex, we use flow matching similarly to [22], where $p_{0}$ is a Dirichlet distribution across all clusters. The model predictions are projected on the simplex using a Von Neumann projection [24]. Features are generated using the process introduced in [20], i.e., $p_{0}$ is the Gaussian distribution, and the feature of the parent cluster is given as conditioning during generation using a FiLM layer [25]. More details can be found in the supplementary material.
@@ 3.6 Minibatch OT-coupling
Minibatch OT-coupling [12, 13] has shown promising results in image generation by encouraging smoother and more linear generative trajectories, which can improve sampling efficiency and model convergence. Instead of sampling independently from the prior and target distributions, OT-coupling aligns samples within each minibatch via an optimal transport plan. Specifically, samples are first drawn from the prior (typically Gaussian), then their corresponding target samples are drawn, and finally, the prior samples are optimally matched (reindexed) to the targets to minimize transport cost. We extend this idea to our setting by applying OT-coupling to the children of an expanded node, treating these children as a minibatch. For each expansion, we solve a local OT problem to align the initial noised samples of children with their corresponding targets.
A crucial constraint is to avoid swapping samples between non-equivalent nodes—i.e., only nodes that are structurally and feature-wise indistinguishable up to noise can be swapped. Swapping non-equivalent nodes would alter the learned data distribution and degrade model fidelity. To keep computational costs manageable, we restrict OT-coupling to groups of children within a single expanded cluster, satisfying the equivalence condition. Since our expansion step produces two children per cluster, OT-coupling simplifies to evaluating just two permutations per cluster, making it extremely lightweight and easily parallelizable using standard tensor operations.
We summarize our theoretical guarantee as follows:
@@
Minibatch OT-coupling in our framework preserves the target distribution while improving the straightness and consistency of the learned generative flow.

### v2

@@ 3.3 Budgeted Expansion and Refinement
Once multi-scale representations for the dataset of hypergraphs are produced by the coarsening procedure, the objective is to train a model able to reverse it. In this section, we define the inverse of the coarsening procedure, which the model will learn to imitate. It comprises two stages: expansion (“upsampling” by duplicating nodes and hyperedges) and refinement (adjusting connectivity, budgets, and features). This process exclusively works on the bipartite representation.
Since our framework only conditions on the total number of nodes in the final hypergraph, we maintain a single node budget vector $\mathbf{b}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ for the node (left-side) partition and discard the hyperedge (right-side) budget. To undo mergings, clusters are recursively split into multiple child nodes. Each child initially copies the parent’s connections and inherits its feature and budget, before a subsequent refinement step, then: i) redistributes the budget among the children, ii) updates their features, and iii) removes edges that should not persist at the finer scale.
@@
Given a bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$ and two expansion size vectors $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$, $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$, denoting the number of duplication for nodes (left side) and hyperedges (right side), respectively. Let $\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b}^{\text{expanded}},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$ denote the expansion of $B$, whose node sets, budgets, and features are given by:
[…]
@@
Expansion thus acts as a clone‑and‑rewire operation: vertices and hyperedges are duplicated, and each child inherits every incident edge of its parent.
@@
Given a bipartite graph $\tilde{B}=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$, an edge selection vector $\mathbf{e}\in\{0,1\}^{|\tilde{\mathcal{E}}|}$, a budget split vector $\mathbf{f}\in[0,1]^{|\tilde{\mathcal{V}}_{L}|}$, satisfying $\sum_{v\in\mathcal{V}_{L}^{(p)}}\mathbf{f}[v]=1$ for all cluster $\mathcal{V}_{L}^{(p)}$ in $\tilde{\mathcal{V}}_{L}$, and two feature refinement vectors $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ with the same dimensions as $\mathbf{F}_{L}^{\text{expanded}}$ and $\mathbf{F}_{R}^{\text{expanded}}$, let $B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\mathcal{E},\lceil\mathbf{b}\odot\mathbf{f}\rfloor,\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ denote the refinement of $\tilde{B}$, where $\mathcal{E}\subseteq\tilde{\mathcal{E}}$ such that the $i$-th edge $e_{(i)}\in\mathcal{E}$ if and only if $\mathbf{e}[i]=1$.
@@
3.4 Probabilistic Modeling
We now present a formalization of our learning framework, generalizing [9, 10]. Let $\{H^{(1)},\ldots,H^{(N)}\}$ denote a set of i.i.d. hypergraph instances. Our objective is to approximate the unknown generative process by learning a distribution $p(H)$. We model the marginal likelihood of each hypergraph $H$ as a sum over the likelihoods of its bipartite representation’s expansion sequences $p(H)=p(B)=\sum_{\varpi\in\Pi(B)}p(\varpi)$, where $\Pi(B)$ denotes the set of valid expansion sequences from a minimal bipartite graph to the full bipartite representation $B$ corresponding to $H$. Each intermediate $B^{(l-1)}$ is generated by expanding and refining its predecessor, in accordance with Definitions 4 and 6.
To simplify notations, we drop the superscript for $\mathbf{F}^{\text{refine}}$ and simply write $\mathbf{F}$. Assuming a Markovian generative structure, the likelihood of a specific expansion sequence $\varpi$ is factorized as:
$p(\varpi)=\overbrace{p(B^{(L)})}^{1}\prod_{l=L}^{1}p(B^{(l-1)}|B^{(l)})=\prod_{l=L}^{1}p(\mathbf{e}^{(l-1)},\mathbf{f}^{(l-1)},\mathbf{F}_{L}^{(l-1)},\mathbf{F}_{R}^{(l-1)}|\tilde{B}^{(l-1)})p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)}).$
To simplify the modeling process and avoid learning two separate distributions $p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})$ and $p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})$, we rearrange terms as follows:
$p(\varpi)=p(\mathbf{e}^{(0)},\mathbf{f}^{(0)},\mathbf{F}_{L}^{(0)},\mathbf{F}_{R}^{(0)}|\tilde{B}^{0})p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})\left[\prod_{l=L-1}^{1}p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})\right],$
where $p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})=p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)}|B^{(L)})$.
We assume that the variables $\mathbf{v}_{L}^{(l)}$ and $\mathbf{v}_{R}^{(l)}$ are conditionally independent of $\tilde{B}^{(l)}$ when conditioned on $B^{(l)}$:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)},\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)}).$ (6)
This allows us to write the combined likelihood as:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)}),$ (7)
which corresponds to the combined expansion and refinement step that inverts coarsening at depth $l$.
During training, the model approximates the right-hand side of (7), learning to generate, for each expansion and refinement step, $\mathbf{v}_{L}$, $\mathbf{v}_{R}$, $\mathbf{e}$, $\mathbf{f}$, and $\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}$. Once trained, the model generates a hypergraph with $N$ nodes through successive expansion and refinement steps:
@@ 3.4 Probabilistic Modeling
    Initialization. Start from a minimal bipartite graph: $B^{(L)}=(\{1\},\{2\},\{(1,2)\},(N))$, consisting of a single node on each side connected by one edge. The left-side node is assigned the full node budget, i.e., $\mathbf{b}=[N]$. If node and hyperedge features need to be generated, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$ are initialized as zero matrices.
@@
    Expansion and refinement. Iteratively expand and refine the current bipartite representation to add details until the desired size is attained: $B^{(l)}\xrightarrow{\text{expand}}\tilde{B}^{(l-1)}\xrightarrow{\text{refine}}B^{(l-1)}$.
@@ 3.5 Minibatch OT-Coupling
Graphs and hypergraphs lack a natural node ordering, making it challenging to align predictions with their targets. After an expansion, the model may produce a permutation of the correct predictions; yet, without alignment, the computed loss will be large despite the prediction being correct. This mismatch introduces significant noise into the learning signal. To address this, we generalize minibatch OT-coupling [12, 13] from image generation to our setting. In its original form, OT-coupling aligns samples from the prior and target distributions within a minibatch via an optimal transport plan, reindexing prior samples to minimize the matching cost. Formally, given prior samples $\mathbf{X}\in\mathbb{R}^{B\times d}$ and targets $\mathbf{Y}\in\mathbb{R}^{B\times d}$, OT-coupling finds:
$\mathbf{P}^{\ast}=\operatorname*{\operatorname{arg\,min}}_{\mathbf{P}\in\Pi_{B}}\|\mathbf{X}\mathbf{P}-\mathbf{Y}\|^{2},$ (8)
where $\mathbf{P}$ is a permutation matrix and $\Pi_{B}$ denotes the set of all $B\times B$ permutations.
In our context, for a given bipartite graph with $n$ nodes, adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$, expanded budgets $\mathbf{b}\in\mathbb{N}^{n}$ and expanded features $\mathbf{F}\in\mathbb{R}^{n\times d}$ where $d$ is the dimension of node or hyperedge features, $\mathbf{X}$ and $\mathbf{Y}$ correspond to prior samples and targets, respectively, and $\mathbf{P}$ is implemented as a node-swapping operation. To avoid bias, swaps are only allowed between equivalent nodes—those structurally and feature-wise indistinguishable up to noise. Swapping non-equivalent nodes would alter the learned distribution and degrade model performance. This leads to the following problem:
$\mathbf{P}^{\ast}=\operatorname*{\operatorname{arg\,min}}_{\mathbf{P}\in\Pi_{n}}\|\mathbf{P}\mathbf{X}-\mathbf{Y}\|^{2}\quad\text{s.t.}\quad\mathbf{P}^{\top}\mathbf{A}\mathbf{P}=\mathbf{A},\;\mathbf{P}\mathbf{b}=\mathbf{b},\;\mathbf{P}\mathbf{F}=\mathbf{F}.$ (9)
To reduce overhead, we only consider permuting children from a single cluster expansion, where equivalence holds naturally. In our implementation, with exactly two or three children per cluster, only two or six permutations are possible, making the operation lightweight and easily parallelizable with standard tensor operations.
@@
Minibatch OT-coupling in our framework preserves the target distribution and produces shorter target paths.
3.6 Generalization to Graphs
Our framework can be directly extended to generate standard graphs by adapting the coarsen–expand procedure of [9]. As in the hypergraph setting, each node begins with a budget of one, which is summed when nodes are merged, while features are aggregated through a weighted average with weights proportional to node budgets. During expansion, clusters propagate their budgets and features to child nodes, and the model predicts how to split the budget among them while refining their features. Graph generation, therefore, starts from a single node with a budget equal to the desired size and proceeds through successive expansion and refinement steps. To ensure stable training and alignment between predicted and target structures, minibatch OT-coupling is applied to groups of expanded nodes exactly as in the hypergraph case.